Estimate the pKa using Pauling's rules for the formation step of the following bases: 1) HPO42-. 2) ClO3-. 3) SeO4-.
Answer #1

According to Paulings Rule,

pKa = 8 - 5p , where p=no.of oxo groups

1]. HPO42- is conjugate base of H2PO4-,

Give them in form of (OH)qP(O)p

For H2PO4-,(OH)2P(O)2,no of oxo groups=2

pKa=8-(5×2)=-2

2].ClO3-  is conjugate base of HClO3,

For HClO3=(OH)Cl(O)2,no of oxo groups=2

pKa=8-(5×2)=-2

3]. SeO42- is conjugate base of HSeO4-

For this,(OH)Se(O)3,no of oxo groups=3

pKa=8-(5×3)=-7